Status Report

ISS On-Orbit Status 10 June 2002

By SpaceRef Editor
June 10, 2002
Filed under , ,

All ISS systems continue to function nominally,
except as noted previously or below.  It is Day 187 for the
Expedition 4 crew (185 days on board ISS).

This morning, the SSRMS (space station remote manipulator system),
steered by FE-1 Peggy Whitson and MS3 Carl Walz, successfully
installed the MBS (mobile base system) mechanically on the MT (mobile
transporter), located at the S0 truss.  After the MBS’s
electromechanical capture claw (MTCL) grappled a capture bar on the
MT at 8:50am EDT, the mobile base was settled down on the MT at
9:03am through the action of the latches and alignment guides.
 [Last night, the MBS was maneuvered by the SSRMS (space
station remote manipulator system) to the pre-install position, about
3 ft above the MT, to allow the two contact surfaces to achieve
temperature equilibrium (max. allowable temperature differential: 50
deg C).  Over night, checkout was performed by CSA on some MBS
subsystems, such as the MBS VDU1 (video distribution unit 1), CLPA
(camera, light & pan/tilt assembly), MTCL powerup and MTCL
checkout open.]

Both crews received kudos from the ground on their great job with
yesterday’s EVA-1, particularly for the careful monitoring of tight
SSRMS clearances during SMDP (SM debris panel) stowage transfer and
APFR (articulated portable foot restraint) egress.  Duration of
the spacewalk was 7h 14m, and all planned tasks were accomplished.
 Only two minor issues were reported: no biomedical data
received for one EVA crewmember, and an initial boot fit problem.
 No changes required to EVA-2.

EVA-2, again by Franklin Chang-Diaz and Philippe Perrin, is scheduled
for tomorrow (to begin approximately 11:00 am EDT).  [The
spacewalk will focus on mating power, data and video umbilicals
between the MBS and MT, including putting inhibits in place for
mating the power cables and conducting tests to verify establishment
of correct power and data paths.  The MBS will then be bolted
down permanently on the MT.  Until the umbilical mating, the
SSRMS provides "keep alive" power to the MBS heaters and
electronics via prime and backup power channels, a capability which
became available after installation of the S0 truss segment during
8A.  A second EVA task is the rearrangement (rotation) of the
POA fixture (payload orbital replacement unit accommodation) on the
MBS, for holding future cargoes as the MBS moves along the truss
railway. Next, Chang-Diaz will hook up the newly configured redundant
cross-wire power jumper cables between MBS and MT, delivered on
STS-111.  Final task will be the relocation of a television
camera from the MBS keel to its top, and the attachment of a bag
containing a contingency MBS extension cable to the MBS.]

The newly introduced O2 prebreathing in the ISS using oxygen from
the Shuttle through 90 ft hoses was so successful that today’s
planned transfer of Shuttle O2 to the ISS Airlock HPGTs
(high-pressure gas tanks) was no longer required.  [The
pre-breathing used 12 lbs of O2 from the Shuttle tanks and only 6 lbs
from the Airlock HPGT (for EMU recharge), less than predicted.
 This does not leave sufficient ullage (free head room) in the
HPGT to justify a transfer.  Nitrogen gas (N2) has already been
transferred (18.9 lbs).]

MS2 Chang-Diaz and MS1 Perrin today performed EMU (extravehicular
mobility unit) maintenance for tomorrow’s EVA, including removing
LiOH (lithium hydroxide) canisters from the EMUs, installing Metox
canisters, replacing batteries and installing used batteries in the
REBA (rechargeable EVA battery assembly).  [Since both
crewmembers had movement or partial movement of the SAFER (simplified
aid for EVA rescue) manual isolation valve during EVA-1, EVA
specialists have recommended adding a step to check the isolation
valves during the EVA egress procedure for both EVA-2 and EVA-3.]

MPLM transfers continue to go very well, with the additional crew
time "bought back" from the cancelled O2 transfer made
available to these activities.  All middeck items, except for
six, have already been transferred.  Updated transfer lists are
uplinked regularly.  [As of last night, 73% of MPLM items
were transferred. This includes all cargoes of five logistics racks,
95% of three logistics racks, 40% of four logistics racks, and 0% of
one logistics rack.]

Crew handovers are proceeding on schedule, but time on  the
crowded UF-2 schedule is precious. To compensate for lacking time,
parallel work time has been introduced to replace dedicated handover
time to some extent.  As of this morning, handover already
amounts to 3h 40m for the CDRs (Onufrienko/Korzun), a bit over 3 h
for FE-1 (Walz/Whitson) and no time yet for FE-2

Yuri Onufrienko and Valery Korzun, assisted by Dan Bursch and Sergei
Treschev, were scheduled for a science experiment with the
Russian/Japanese HDTV (high definition TV) experiment.  The
session involves taking high-definition video of the subjects during
"interviews", showing their facial features for medical

Dan Bursch completed the long-planned swap-out of four laptops,
required to refresh CMOS batteries in the machines.  Preparatory
to UF-2, Carl Walz had completed a task-listed laptop rearrangement
procedure at an earlier time, swapping out two PCS (personal computer
system) laptops and two SSC (space station computer) laptops.
 This had to be done prior to UF-2 since it positioned the
laptops so that only four laptop swaps were required today, rather
than six.  [The rearrangement exchanged cards & hard
drives of the SSC "FGB" router 1 and PCS "SM"
laptops, and of the SSC 5 and PCS "Lab RWS" laptops.]

CDR Onufrienko, monitored by Korzun, terminated the regeneration
cycle on absorbent bed #2 of the Russian BMP micropurification unit
and switched it to Purify mode.  Both BMP channels are now again
in this mode, until the next regeneration of the recyclable

MS1 Philippe Perrin, joined by CDR "Taco" Cockrell, had an
live TV event with French media at 11:27 am EDT.  Although
located in the Lab during the exchange, they used Shuttle comm
assets, with the Shuttle camcorder.   

The formal Change-of-Command ceremony between CDR Yuri Onufrienko and
CDR Valery Korzun, originally planned for 2:12 pm EDT, was delayed by
a false fire alarm from the FGB (the smoke sensor appears to have
been triggered by excessive dust in the air stirred up by logistics
transfer operations). The ceremony was to be downlinked via station
assets with Ku-band/video and S-band/audio coverage.  This will
make the rotation of the Expedition 4 crew with the new station
residents of Expedition 5 "official".

Todayâs middeck payload checks were spread around between Carl Walz,
Dan Bursch and "Paco" Lockhart, while Sergei Treschev
performed the regular daily ISS UF-2 payload status check, using the
computerized procedures viewer (OSTPV) as reference.

After the hard failure of CMG1 (control moment gyro #1) on Saturday,
6/8, attitude control of the "stack" using three CMGs
(control moment gyros) in the yaw-biased attitude, selected to keep
the MBS out of shadow before installation, has been going very well,
with more margin than predicted.  However, in order to assure
minimal stress on the remaining CMGs, the Shuttle will be used for
attitude maneuvers during the docked phase.  Attitude maneuvers
and attitude hold periods after Shuttle departure will be performed
by the Russian MCS (motion control system) thrusters.  The CMGs
will continue to be used for momentum management, after maneuver
completions.  [All data currently available indicate a
failure of the spin motor bearing. NASA/JSC has formed three special
teams to work on the CMG issue: one for investigating the root cause,
one for looking at near-term operations and options with three CMGs,
and the third to work on the logistics of readying a CMG spare unit
and the actual R&R (removal and replacement) of the unit in

Last night the MCOR (medium-rate communications outage recorder)
was powered down (again) to prevent it from beeping when reaching its
operational limit of 46 deg C.  The crew heard its beeps this
morning from the Lab when MCOR was powered up.

Later today, at about 4:53 pm EDT, the Shuttle will perform the first
of three planned reboosts of the stack.  The one-hour maneuver
is automatic, controlled by the Endeavour’s DAP (digital autopilot)
with special software to assure thruster firings are timed to avoid
structural excitation (vibrational resonance).  Expected
increase in mean altitude is about 1 naut. mi. (1.852 km).
 [The idea of a fourth reboost has been discarded.  It
was intended to make up for the propellants to be used by Progress 8P
for its test flyaround on 6/28.  MCC-Moscow is now looking into
conducting the systems test with minimum propellant expenditure.]

Today’s target areas for the combined Shuttle/ISS earth
observation programs were Angola, Zaire (burning of
agricultural fields is well under way in the region. Crew was to
document fires/point sources of smoke, as well as regional smoke
palls. Normally exposed frames permit estimates of smoke thickness,
whereas slightly overexposed views reveal the maximum extent of smoke
, Mesopotamian Marshlands, Iraq (where water
resources have long been unevenly distributed, recent political
actions are further complicating the situation. ISS tracked over what
were once lush wetlands at the head of the Persian Gulf, and the crew
was asked to  photograph the remaining patches of marshlands
[dark, vegetated areas], perhaps taking advantage of sun glint off
the canals and small ponds to record any surface water in the region.
They were to note any canal and/or construction)
, Baikal and
Hovsgol Rifts, Potential New Rift
(crew to photograph the
north-trending valleys that lie west of the Lake Hovsgol rift valley.
Newly returned photos from STS 110 have revealed what may be a third
rift W of Hovsgol. Nadir views of the transform fault zone that links
Hovsgol with the great Baikal rift are needed as well.  Crew was
asked to look right as they crossed southern Lake Baikal to record
volcanic centers at the SW corner; any plumes of air or water
pollution from cellulose plants along the south shore are also of
, Caspian Sea, Volga and Ural River Deltas
(over the past two decades Caspian Sea level has been rising, and
the impact has been greatest on the low-relief parts of the coast.
Effects have been notable on the deltas of the Volga and Ural Rivers,
Kazakhstan, where grazing and farm lands have been inundated. Levees
along dredged channels in the outer deltas have been breached and
partially destroyed.  Crew was asked to look left of track and
document early summer water levels, as well as vestiges of
, Southern and Central France (a
substantial blanket of clouds, which is drawing moisture from both
the Atlantic and the Mediterranean, lies over France. It was unlikely
that conditions will clear, but if that occurs, patches of the
western Pyrenees and upper Rhone rift valley might become
, Amazon River Delta (conditions appeared
favorable for photographing the mouths of the Amazon River, with its
many islands and channels. Of interest: looking left and recording
the condition of the wetlands, as well as any sediment plumes being
carried out to sea. A thick apron of mud-laden water commonly lies
adjacent to the delta, and photos of the seaward extent of the goo
are needed as well)
, Peruvian Mountain Glaciers (small
mountain glaciers cap several peaks of the Peruvian High Andes
 the crew could see them both left and right of track.
Glaciologists and climate modelers are monitoring the extent of such
ice masses at equatorial latitudes; many are rapidly retreating as a
result of short- and/or long-term climate changes)
, Soufriere
Hills Volcano, Montserrat
(Montserrat with its active volcano
was immediately left of track, and clouds were moving out of the
area. Conditions should have been be favorable for near-nadir views
of the mountains, new lava flows and ash deposits, and any plumes of
ash or steam)
, North Coast of France (thick clouds will
likely remain over the region. Should conditions clear the crew may
have been able to see Calais and patches of northern France)
Costa Rica-Nicaragua Border (clouds appeared to be moving
away from the west coast of Central America but weather conditions
are unsettled. The crew may have had a view of San Cristobal volcano,
which has been sending incandescent ash into the air, and Managua,
where there has been recent flooding. If the crew had clear
conditions, photos of both were needed. Lago Nicaragua and northern
Costa Rica were immediately right of track)
, Tropical Storm
Boris, Fires, Colima Volcano
(ISS passed directly over
Tropical Storm Boris as it approached the Mexican coast; it is moving
slowly WNW-ward and, as yet, has not attained hurricane force.
Volcano Colima  has sent new lava flows down the W and SW flanks
of the mountain and incandescent lava avalanches have been common
during past few days. Of interest: documenting any plumes, as well as
flows and ash deposits. Wildfires in the sierras south of Puerto
Vallarta are still uncontained, crew was to record the extent of both
fires and burn scars)
, Baja California, Sonora, Arizona,
Colorado Fires
(from the delta of the Colorado River in
northern Baja California northward across Sonora, Arizona and
Colorado, wildfires have been rampant in recent days. Crew was asked
to document these summer fires in a region that has been in a dry
cycle for a couple of years)
, SE Australia, Warrega River
(as the ISS traveled NE up the Darling River drainage,
crew was to look left to record the Warrega River and the broad
megafan that it is constructing.  They also were to look right
as well to photograph megafans on more southerly tributaries of the
,  and Great Barrier Reef, Coral Bleaching
(bleaching of corals on the Great Barrier Reef, in response to
warming water temperatures, has been reported in the region. Crew was
to take nadir mapping views of the southern Reef complex as ISS
crossed. The crew also had an excellent opportunity to look left and
photograph the whole barrier reef).

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 6:00 am

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode). Vozdukh
CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed).
 U.S. CDRA CO2 scrubber is Off.  BMP Harmful Impurities
unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

SM Working Compartment:  Pressure (mmHg) — 753,
temperature (deg C) — 26.7, ppO2 (mmHg) — 144.7, ppCO2 (mmHg) —
SM Transfer Compartment:  Pressure (mmHg) — 752,
temperature (deg C) — 20.3.
FGB Cabin:  Pressure (mmHg) — 744, temperature (deg C)
— 21.3.
Node:  Pressure (mmHg) — 735.62, temperature (deg C) —
22.8 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 737.64, temperature (deg C)
— 23.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 737.54,
temperature (deg C) — 22.9; shell heater temp (deg C) — 22.6, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 22.3
PMA-2:  Shell heater temp (deg C) — 20.0

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) have been removed for return to

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in
Autotrack mode.
SM batteries: (data not available)
FGB: Battery #3 is offline (ROM mode). All other batteries (5)
are in ãpartial chargeä mode.
Plasma Contactor Unit PCU-1 in Discharge mode; PCU-2 in Discharge

Thermal Control Systems:
Air conditioner SKV-1 is Off. SKV-2 is On.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM on Standby.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating

SSRMS/Canadarm2 at MBS.
RWS (robotics workstations) are On.
SSRMS Prime string Wrist Roll (WR) joint to be replaced on UF-2

ISS Orbit (as of this morning, 6:08 am EDT [=
Mean altitude — 385.2 km
Apogee — 389.1 km
Perigee — 381.3 km
Period — 92.2 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0005744
Orbits per 24-hr. day — 15.60
Altitude decrease — 150 m (mean) in last 24 hours
Revolutions since FGB/Zarya launch (Nov. ’98) — 20297
Current Flight Attitude — LVLH (local vertical/local horizontal =
“earth-fixed’: z-axis in local vertical, x-axis in velocity vector
[yaw: 5 deg, pitch: 23 deg., roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times,

SpaceRef staff editor.